Moskowskaja, 267, Brest 224017 Belarus Brest State Technical University Rector: Petr Poita

10 April 2006 Deputy Ambassador Mr. Takayuki KOIKE

Brest State Technical University is going to held the ICNNAI-2006 conference and asking Japanese Embassy to be one of the co-sponsors.

Beside this application form, we attached (1) Confirmation; (2) Call for Papers; (3) Financial plan; and (4) About the group who hold the conference.